OpenStudy (anonymous):

Solve for h: A=1/2 b(h+y)

OpenStudy (anonymous):

\[ A =\frac 1 2 b(h+y) \] Is this your problem?

OpenStudy (anonymous):

\[ 2 A= bh +by\\ bh = 2 A -by\\ h =\frac {2 A -by}b=\frac {2A}b -y \]

OpenStudy (anonymous):

wow thank you so much!